Due to a number of lawsuits related to toxic wastes, a major chemical manufacturer has recently experienced a market reevalua\

$1,273 | $1,000 | $7,783 | $550 | $450

please show work in written form (not excel) thank you.
Answer #1

price of coupon = Coupon payment per period * [1-(1+i)^-n]/i + par value/(1+i)^n

i = interest rate per period = 0.16/2

n = number of periods = 15*2 = 30

coupon rate per period = 1000*8%/2 = 80/2

=>

Current value = 80/2 * [1-(1+0.16/2)^-30]/(0.16/2) + 1000/(1+0.16/2)^-30

= 549.68

= 550 ...........(ans)